OpenStudy (anonymous):

3x – 7y = 5 9y = 5x + 5 A. (–3, 2) B. (8, 5) C. (–10, –5) D. (0, 0)

OpenStudy (shamil98):

3x - 7y = 5 9y = 5x + 5 3x - 7y = 5 y = 5/ 9x + 5/9 3x - 7 (5/9 + 5/9) = 5

OpenStudy (anonymous):

so it would be b?

OpenStudy (shamil98):

well plug in the values to equation is another option. 3(8) - 7(5) = 5 24 - 35 does not equal 5 so no.

OpenStudy (anonymous):

okkaay. then which one? .-.

OpenStudy (shamil98):

Plug the values in!, You can cross out option D , because 3(0) - 8(0) does not equal 5.

OpenStudy (anonymous):

A?

OpenStudy (shamil98):

It is C, please try and do the problem rather than guessing next time.

OpenStudy (anonymous):

ok